According to the National Center for Health Statistics, the mean height for adult women is μ=63.8 inches, with a standard deviation of σ=2.8 inches. Find the proportion of adult women taller than 5’4”.

Write as a decimal and round to the nearest hundredth.
Answer #1

Consider, X= height for adult women

ΧN Nμ = 63.8, σ = 2.8)

The proportion of adult women taller than 5’4”.

P[X>64]......................................................................5’4” = 64 inch

64 - 63.8 = P[Z> 2.8

=P[Z>0.0714]

=1-PZ < 0.0714

=1-0.5285............................by using Z table or by using Excel command =NORMSDIST(0.0714)

=0.4715

Therefore,  The proportion of adult women taller than 5’4” = 0.47